### v1

# SVG360: Multi-View SVG Generation with Geometric and Color Consistency from a Single SVG

Abstract
Scalable Vector Graphics (SVGs) are central to modern design workflows, offering scaling without distortion and precise editability. However, for single object SVGs, generating multi-view consistent SVGs from a single-view input remains underexplored. We present a three stage framework that produces multi-view SVGs with geometric and color consistency from a single SVG input. First, the rasterized input is lifted to a 3D representation and rendered under target camera poses, producing multi-view images of the object. Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining. Finally, during the raster to vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics. The resulting SVGs exhibit strong geometric and color consistency across views, significantly reduce redundant paths, and retain fine structural details. This work bridges generative modeling and structured vector representation, providing a scalable route to single input, object level multi-view SVG generation and supporting applications such as asset creation and semantic vector editing.
1 Introduction
Editable vector graphics are central to modern design workflows, providing resolution independence, structural clarity, and precise editability. However, generating multi view consistent and fully editable vector graphics for a given object remains an open challenge. When designers attempt to present an existing vector illustration from a new viewpoint, they often need to manually redraw or adjust its structure to approximate the desired perspective. This manual process is both time consuming and prone to geometric and stylistic inconsistencies. Adobe Illustrator’s Turntable (Beta) [1] is one of the few publicly available end-to-end tools that generate multiple views from a single 2D vector artwork, yet it provides constrained angle control and discloses no technical specification.
The task of automatically generating multi-view consistent and fully editable SVGs, however, has not been extensively explored. We define this task as synthesizing a set of path level editable SVGs that preserve geometric and stylistic coherence across viewpoints. Solving this problem is particularly valuable for creative workflows such as turntable style visualization, multi-view logo and icon generation, and geometry consistent SVG asset editing. Bridging the gap between generative modeling and structured vector representation is therefore critical for enabling scalable and reliable design pipelines.
Recent advances in generative AI, particularly diffusion models[10, 29, 25, 30, 26, 46, 23], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [20, 33, 21, 38].
Yet, these outputs remain raster based and lack the structured semantics and fine grained editability inherent to vector graphics. Meanwhile, existing vectorization approaches such as DiffVG [18], DeepSVG [3], and Im2Vec [28] focus on single view reconstruction and fail to maintain geometric consistency across views. This gap motivates a core objective: enabling generative models to produce multi view coherent and structurally editable representations directly in the SVG domain.
To realize this objective, we propose a unified generative to vector framework that produces multi-view, geometrically consistent, fully editable SVGs from a single input. Our method consists of three stages. It first performs three dimensional aware multi-view generation to obtain geometry plausible novel views. Then it performs cross-view refinement through spatially aligned segmentation and appearance harmonization, enhancing the consistency of object contours and colors across views in the SVG style domain. Finally, leveraging the multi-view consistent segmentation results, we perform raster-to-vector conversion followed by path consolidation and color correction, yielding compact, coherent, and easily editable multi-view SVGs.
On multi-view SVG generation, our method produces noticeably cleaner and more stable results than existing vectorization tools. Compared with Adobe Illustrator’s Turntable, the path counts of our generated views stay much closer to the structure of the input SVG, reducing the path-count deviation from the input SVG by 26.5%. The variation in color usage between adjacent viewpoints is also reduced by 83.5%, indicating a far more consistent preservation of the input color design across rotations. In addition, our method lowers the average number of paths by 11.6%, resulting in lighter and more editable multi-view SVG sequences. The generated views exhibit coherent appearance transitions and show strong potential for creative design workflows, turntable-style visualization, and semantically aligned vector editing. Our contributions are summarized as follows:
Unified Generative to Vector Framework. We introduce the first end-to-end framework to directly synthesize multi view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
Spatially-Aware Cross-View Segmentation Refinement. We design a spatially-aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
Structure Preserving Vector Consolidation. We develop a vector domain consolidation strategy that merges redundant paths while maintaining semantic boundaries and structural coherence, yielding compact and semantically stable SVG assets suitable for real world design workflows.
The remainder of this paper is organized as follows: Section 2 reviews related work on multi-view SVG generation and the open gap; Section 3 presents our three stage framework; Section 4 presents experimental comparisons and ablation studies; Section 5 concludes and discusses limitations.
2 Related Work
2.1 Single image to multi view generation
Turning one reference view into a usable turntable sequence has become a key precursor to our goal of editable multi-view SVGs. Existing approaches fall into two major categories:
Diffusion-based multi-view image synthesis builds on high-fidelity 2D diffusion priors. These methods condition the denoising trajectory on target camera poses to synthesize each view directly in the image domain. Representative works include Zero-1-to-3 [20] and Zero123++ [33], where the former targets novel-view images rather than explicit 3D, and the latter improves cross-view consistency through better conditioning and training. Further, SyncDreamer [21] explicitly couples views during sampling via synchronized diffusion to align geometry and color, and SV3D [38] leverages video-latent diffusion to generate orbital view sequences from a single image. Despite strong appearance fidelity, residual independence across views often leads to geometric and color drift on thin structures and self occlusions.
3D reconstruction and generation from single images. These methods predict a 3D asset from one or a few views and render all novel viewpoints from the shared geometry, which naturally improves multi-view geometric consistency over 2D diffusion synthesis. Representative approaches include mesh generators such as GET3D, PolyGen, MeshDiffusion [8, 24, 22], single-image reconstruction methods like LRM, TripoSR, UniQue3D [11, 36, 40], and latent 3D models including Shape-E and Trellis [14, 41]. However, their outputs remain rasterized images without explicit vector paths or cross-view path correspondences, making them unsuitable for SVG-oriented workflows.
3D Gaussian Splatting (3DGS) as a shared representation. 3DGS [15] represents a scene with anisotropic Gaussians and renders by fast differentiable splatting, so all views reproject the same underlying 3D field. This often yields strong cross view geometric stability and efficient multi view rendering. Recent works combine diffusion priors with 3DGS for rapid 3D asset creation [35, 7], and single image variants enable ultra fast inference [34]. Although 3DGS provides efficient and geometrically consistent multi-view rendering, its outputs remain raster-based and do not explicitly expose vector path topology. Consequently, converting rendered contours into clean, closed, and editable paths requires an additional vectorization stage.
2.2 Raster to Vector Generation
Transforming multi-view consistent rasters into structured vector representations remains challenging because recovering continuous topology and geometry from discrete pixels is inherently ill-posed. Classical contour tracing methods such as Potrace and AutoTrace detect boundaries and fit Bézier curves or polygons. They perform well on icons or low-color imagery, but on smooth gradients and complex textures they often produce jagged, fragmented, or redundant paths.
Learning-based approaches model vector paths parametrically. DiffVG [18] introduces a differentiable renderer for gradient-based optimization, and DeepSVG [3] and Im2Vec [28] leverage Transformer or VAE designs to improve geometric continuity and structural controllability. LIVE [45] decomposes images into hierarchical layers to preserve global topology, yet can over-segment fine details. More recent methods such as VectorFusion [13] and StarVector [39] incorporate diffusion priors or region fusion to obtain smoother and more coherent paths.
Despite these advances, all vectorization methods still operate directly on pixel-level color transitions, which makes complex objects prone to fragmented or redundant paths. Introducing semantic segmentation before vectorization greatly simplifies the problem by isolating coherent part regions, enabling cleaner per-part conversion and more stable geometric structures. This highlights the need for a semantic layer, especially when extending vectorization to multi-view settings.
2.3 Segmentation and Cross View Consistency
While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation raises an additional requirement: achieving more consistent part assignments across viewpoints greatly improves cross-view coherence and visual stability. Yet conventional segmentation relies mainly on per-image pixel cues, making such consistency difficult; a spatially stable segmentation step is therefore needed before vectorization to provide more reliable part structures across views.
Early two-dimensional methods [9, 4, 42, 17] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection. Foundation-scale models such as SAM, HQ-SAM, and SAM2 [16, 5, 27] provide accurate boundaries and promptable interaction, thus serving as strong building blocks for multi-view scenarios. However, their propagation mechanisms are primarily temporal rather than spatial, which limits their ability to maintain consistent part correspondences across viewpoints in static 3D scenes.
Inspired by video object segmentation, memory-based propagation [6, 43] maintains features across frames to keep masks consistent over time. By analogy, multi-view consistency can be formulated as a spatial sequence problem on the viewing sphere, where each camera view is treated as a pseudo-temporal step and neighboring views act as spatial references.
Existing work either stabilizes masks in raster space or improves single-view vectorization quality, but a bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored. This gap motivates our approach, which first achieves part-aware, cross-view consistent segmentation and then leverages these semantic units to drive compact and editable vector representations across views.
3 Method
Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis. A lightweight LoRA [12] tuned FLUX [2] model is then applied to harmonize their appearance. The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere. During segmentation propagation, the Spatial Memory Selector retrieves the geometrically most relevant memory entries. For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints. The resulting masks are vectorized by VTracer [37] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
As illustrated in Figure 1, our SVG360 framework begins with a 3D-aware rasterization stage (§3.1) that lifts a single-view SVG into geometrically plausible multi-view rasters. We then perform spatially aligned segmentation propagation3.2) to establish part-level correspondences across multi-view. Finally, the vector consolidation stage (§3.3) converts the segmented rasters into compact, fully editable multi-view SVGs with consistent geometry and color.
3.1 Multi-View Raster Generation and Harmonization
Multi-view raster generation. Generating accurate and geometrically consistent multi-view images from a single SVG input is challenging, as purely 2D generation approaches struggle to ensure structural coherence across views. To address this, we adopt a generative 3D-based approach to synthesize multi-view images.
Although several recent single-image 3D methods (e.g., Sparc3D [19] and Hi3DGen [44]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs. Such geometry-only representations are unsuitable for our color consistent raster to SVG multi-view workflow. Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [41] as the backbone 3D generator. Trellis offers a balanced trade-off between geometric accuracy and texture consistency, and remains efficient and practical for producing coherent multi-view rasters from a single SVG input.
Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [41]. Its mesh decoder builds on FlexiCubes [32], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface. While both decoders yield high-fidelity results, converting the latent to a mesh and then rasterizing from many viewpoints introduces extra discretization and baking steps, which may accumulate appearance deviations relative to the input SVG style and incur nontrivial runtime overhead. For efficiency and fidelity, we directly use the 3D Gaussian decoder and render dense multi-view rasters via Gaussian splatting, which preserves the appearance statistics needed by our downstream segmentation and vectorization stages.
Raster harmonization.
While a 3D proxy preserves global geometric consistency across viewpoints, the reconstruction can still exhibit local errors, especially along edges and curved surfaces. View dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion. We apply appearance harmonization using FLUX.1-dev [2] fine-tuned with a LoRA [12] trained on flat-design SVG data. The harmonized renders increase intra-region homogeneity and boundary contrast, suppress spurious shading edges, and reduce color variation across views. The resulting rasters are both view-consistent and faithful to SVG-style appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.
3.2 Spatially Aligned Segmentation Propagation
After appearance harmonization, the multi-view rasters exhibit more uniform regions and sharper boundaries. To obtain cleaner and semantically stable vector paths in the subsequent vectorization stage, we need to maintain consistency of part-level identity across views. Relying on single view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence. Multi-view SVG generation inherently requires segmentation to be continuous in space. Accordingly, we spatialize the continuous segmentation paradigm of SAM2 by replacing temporal adjacency with spatial neighborhood on the viewing sphere and propagating masks along that sphere. Each target view is segmented under guidance from its nearest spatial neighbors, thereby maintaining part identity and boundary consistency at a global level.
Spatially Sequential Scheduling. We uniformly sample camera viewpoints on a unit sphere to obtain a globally distributed set of directions covering the entire view space. The sampling density and angular step are adjustable, allowing flexible control over the number of viewpoints and their spatial distribution according to task requirements. This configuration achieves globally balanced coverage even with a limited number of samples, forming a stable spatial adjacency structure that serves as the geometric foundation for subsequent cross-view propagation and visibility-aware reprojection. Although the sampled views have no intrinsic temporal order, a pseudo-sequential traversal helps maintain continuity during segmentation propagation. Starting from the front view aligned with the input SVG (denoted as $\theta_{0}$), each subsequent view is chosen as the unvisited one with the smallest angular distance from the current view, until all views are processed:
$\theta_{t+1}=\arg\min_{\theta\in\text{unvisited}}d(\theta_{t},\theta).$ (1)
This nearest-neighbor traversal minimizes geometric discontinuities between adjacent views, producing smoother and more stable mask propagation. To measure proximity between any two viewpoints, we define their angular distance as:
$d(\theta_{i},\theta_{j})=\mathrm{atan2}\big(\|u_{i}\times u_{j}\|,\,\mathrm{clip}(u_{i}\cdot u_{j},\,-1,\,1)\big),$ (2)
where $u(\psi,\phi)$ denotes the unit viewing direction for a camera defined by horizontal rotation (yaw, $\psi$) and vertical rotation (pitch, $\phi$):
$u(\psi,\phi)=[\cos\phi\cos\psi,\;\sin\phi,\;\cos\phi\sin\psi].$ (3)
This $\mathrm{atan2}$-based formulation is numerically more stable than $\arccos$, particularly when two directions are nearly parallel or opposite, and directly corresponds to the great-circle distance on the unit sphere. To further enhance traversal smoothness, we apply a lightweight two-segment swap (2-opt) optimization to remove local discontinuities and maintain continuous transitions between neighboring viewpoints.
Reference View Selection. After determining the traversal order, we construct a reference set for each target view to reuse information from geometrically adjacent and previously processed views. For each target view $\theta$, we select its $k$ nearest processed neighbors within an angular threshold $\tau$, ensuring that the guidance originates from spatially reliable local contexts. We empirically set $k=6$ and $\tau=75^{\circ}$ in all experiments. This localized reference selection, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
Key Frame Initialization and Filtering.
Segmentation starts from the key view $\theta_{0}$, which corresponds to the input SVG. After obtaining initial masks from the automatic segmentation model, we apply a lightweight post-processing to reduce redundancy and improve stability for subsequent propagation. Small regions smaller than either 200 pixels or 0.05% of the image area (whichever is larger) are removed, a single morphological closing with an elliptical kernel sized to about 0.5% of the shorter image dimension is used to smooth boundaries, and overlapping masks are suppressed using non-maximum suppression with an IoU threshold of 0.5. The cleaned masks serve as prompt inputs to the spatially SAM2, providing stable guidance for the following multi-view segmentation process.
Residual Discovery.
Since the initial segmentation relies solely on the key view $\theta_{0}$, previously unseen parts that appear in other viewpoints may remain undetected. After Spatial-SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions. For each processed view $\theta_{t}$, we compute the uncovered foreground ratio $r$ based on the union of all current masks. If $r$ exceeds 5% of the estimated foreground area, the same automatic segmentation procedure used for the key view is reactivated, but restricted to the uncovered regions. Newly discovered masks are then merged into the existing mask set and propagated to subsequent views through Spatial SAM2 to maintain cross-view consistency. The overall process is summarized in Algorithm 1.
Input: Set of views $\{\theta_{0},\dots,\theta_{N}\}$ with their current segmentation masks
Output: Updated mask set with newly discovered parts
$k\leftarrow 0$; // iteration counter
$T_{\max}\leftarrow 3$; // maximum number of passes
repeat
  for each view $\theta_{t}$ in spatial order do
     Compute uncovered foreground ratio $r$ using the union of all masks in $\theta_{t}$;
     if $r>5\%$ then
        Re-run automatic segmentation (same mode as key view), restricted to uncovered regions;
        Merge new masks into existing results;
        Propagate updates to neighboring views via Spatial-SAM2;
 $k\leftarrow k+1$;
until no new parts are found or $k\geq T_{\max}$;
Algorithm 1 Residual Discovery Loop
3.3 Vector Consolidation
Figure 2: Qualitative comparison. The figure summarizes representative issues observed in Adobe Turntable: (a) geometric inconsistencies across adjacent views; (b) cluttered structures arising from overlapping thin components; (c) merging of parts with similar colors; (d) missing regions in certain viewpoints; (e) color drift and gradual loss of small details across views. Our method produces multi-view SVGs with stable geometry, clear part separation, and consistent color appearance.
After segmentation, we crop each part into a transparent raster and feed it to VTracer [37], which converts pixel regions into path sets by color layering and spline fitting. Processing per part rather than per image reduces cross-part interference and yields a more compact path topology. This simplifies later consolidation and improves cross-view correspondence.
Direct vectorization often produces redundant small fills and micro strokes. We apply a light consolidation in the vector domain: sparsify colors using filled-area statistics to keep dominant colors and fold near-duplicates; clean only stroke-only micro paths to suppress artifacts from anti-aliasing; finally merge all parts within a view into a single SVG. These steps use fixed yet resolution-aware thresholds so the procedure is simple and efficient. This consolidation reduces the number of paths and colors, improves topological simplicity and editability, and yields more consistent boundaries across neighboring views.
To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric [31]. We add a near-black bias to stabilize dark tones: colors in low lightness and low chroma preferentially map to black, otherwise to the closest non-black reference color. This preserves semantic colors while suppressing banding due to shadows and reflections. In practice we preserve alpha values, resolve inherited attributes into explicit colors, and remove empty attributes so that rendering is consistent across editors. Together, color alignment unifies appearance and consolidation compresses structure, improving cross-view consistency and the editability of the final SVGs.
4 Experiments
4.1 Implementation Details
For the 3D generation stage, we employ the TRELLIS-image-large model to generate multi-view rasterizations from the input SVG. To harmonize the rendered appearance, lightweight LoRA adapter is trained to adapt FLUX.1-dev to SVG-style appearance, following standard PEFT settings (rank=32). The text encoder is frozen, and the adapter is trained for 1,000 steps. All experiments are conducted on a single NVIDIA A100-80GB PCIe GPU.
4.2 Evaluation Metrics
We use two metrics to capture cross-view stability. $\mathrm{RMSE}_{\text{path}}$ measures how much each view’s path count deviates from the input SVG. $\overline{\Delta N}_{\text{color,nbr}}$ measures the change in color count between adjacent views, lower indicates more stable color behavior.
4.3 Comparison
Baseline Adobe Illustrator Turntable (Beta) is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork. Its internal implementation, view grid, and technical specifications are not disclosed, and no programmatic interface is provided. We therefore use Turntable as the industrial baseline for our full pipeline comparison.
According to the official documentation, Turntable supports approximately $\pm 120^{\circ}$ horizontal rotation in yaw (sampled every $15^{\circ}$) and three pitch levels at $\{-45^{\circ},0^{\circ},+45^{\circ}\}$. For fair comparison, we adopt the same camera poses defined by the Turntable configuration in our experiments.
Because Turntable only allows manual per-view SVG export, we evaluate this baseline on 10 representative SVG examples from which complete view sequences can be consistently obtained.
Quantitative Comparison Table 1 reports the numerical statistics for Adobe Turntable and our method. Our approach produces fewer paths and colors per view, indicating cleaner vector topology and a more compact palette. For cross-view stability, our method achieves lower $\mathrm{RMSE}_{\text{path}}$ and a substantially lower $\overline{\Delta N}_{\text{color,nbr}}$, showing that the generated views remain closer to the original SVG structure and maintain more consistent color behavior across viewpoints. Together, these results confirm that our multi-view SVGs are structurally simpler and exhibit smoother appearance transitions.
Table 1: Comparison of multi-view SVG statistics between Adobe Turntable and our method.
Metric Adobe Turntable Ours
$\overline{N}_{\text{path}}$ 29.59 26.15
$\overline{N}_{\text{color}}$ 4.80 3.37
$\mathrm{RMSE}_{\text{path}}$  ( $\downarrow$ ) 16.37 12.03
$\overline{\Delta N}_{\text{color,nbr}}$  ( $\downarrow$ ) 2.12 0.35
Qualitative Comparison. Figure 2 summarizes the typical failure patterns we observe when applying Adobe Turntable to single-SVG multi-view generation. Across multiple objects and views, Adobe Turntable frequently exhibits geometric inconsistencies, structural confusion, unstable camera poses, view-dependent clipping, and noticeable color drift. In contrast, our method maintains coherent geometry, clear part separation, and consistent appearance across views.
In terms of geometry, object structures do not remain consistent across views. In example (a), components that should be rigid, such as the roof rack and the headlights—change noticeably between yaw $0^{\circ}$ and $90^{\circ}$, with the rack switching from two crossbars to a flat plate and the paired headlights collapsing into a single central light. Examples (b) and (c) further show that when the input contains multiple thin or closely spaced parts, the generated views often introduce overlaps, clutter, or unintended merging, making individual components difficult to distinguish.
Viewpoint-related inconsistencies are also apparent. In examples (a) and (e), the result at yaw $90^{\circ}$ deviates from a clean side view, exhibiting a slight tilt and geometric deformation. When the input SVG spans a large spatial extent, Turntable does not normalize its scale, causing portions of the object to be clipped outside the rendering canvas, as observed in both (a) and (e).
Color stability across views can also be unreliable. In example (e), the sesame seeds on the burger change brightness across viewpoints, and those on the far side gradually diminish as the camera rotates. In example (c), regions with similar color are merged into a single path, reducing structural separation.
In contrast, our method relies on explicit 3D representation, controlled camera poses, and consistent color and path processing. As a result, the generated multi-view SVGs maintain stable geometry, clear part separation, and coherent appearance across all viewpoints.
4.4 Ablation Study
Figure 3: Segmentation comparison of Spatial-SAM2 and the original SAM2 tracking mode at view 47. The bottom row shows the each six reference views used for memory support, with view 0 serving as the initial key frame. Results shown here correspond to the first iteration before any subsequent refinement.
Figure 3 illustrates the difference between the original Tracking SAM2(Video Segmentation) and our Spatial SAM2 when propagating masks across multi-view settings. The propagation mechanism of Tracking Mode SAM2 is fundamentally linear and one-directional: it progresses strictly along the view sequence and therefore relies on the temporally adjacent set of reference views (views 41–46). While such linear propagation is appropriate in video scenarios, it introduces a notable limitation in static multi-view renderings. If certain structures are not robustly tracked over several consecutive views, the resulting errors accumulate and become amplified along the propagation chain. Consequently, at views with large spatial offsets, the tracking mode frequently fails to recover local structures such as wheels, leading to partial or complete omission.
Our Spatial SAM2, in contrast, does not depend on single-direction propagation. Instead, it selects reference views based on true geometric proximity on the viewing sphere. Besides leveraging temporally adjacent views such as 44–46, it also incorporates views like 24–26, which may be far apart in the temporal order but remain spatially close to the target view. Since reference information is validated jointly across multiple spatial neighbors rather than flowing along a single temporal chain, Spatial SAM2 effectively suppresses error accumulation and enhances segmentation robustness. As a result, even at spatially distant viewpoints, structures such as wheels remain stable and consistently segmented across views.
Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation free VTracer baseline.
As shown in Figure 4, we further compare the impact of four segmentation strategies on the final SVG outputs. Spatial-SAM2 produces the most stable and coherent results. The tracking mode of SAM2 still exhibits occasional misses in certain viewpoints. The Auto mode (i.e., SAM2’s Segment Everything) tends to over-fragment continuous structures due to the lack of cross-view constraints, resulting in inconsistent topology across views. When segmentation is entirely omitted, many parts fail to survive the vectorization stage and cannot be reliably reconstructed, leaving only a few coarse contours. These findings confirm that enforcing spatially consistent segmentation is critical for producing high-quality multi-view SVGs.
5 Conclusion
We presented a unified generative-to-vector framework that produces geometrically consistent and fully editable multi-view SVGs from a single input illustration. By integrating 3D-aware multi-view generation, spatially aligned cross-view refinement, and vector-domain consolidation, our approach bridges raster-based synthesis with structured vector representations and enables practical multi-view SVG editing workflows.
Despite these improvements, several challenges remain. Part-level consistency is still constrained by the limitations of current segmentation models, particularly on objects with intricate topology or numerous fine-grained components. Our framework currently focuses on single-object inputs and does not yet extend to scene-level vector graphics requiring instance reasoning or world-coordinate camera transformations. Furthermore, vectorization remains most reliable for closed, well-formed paths, while open or stroke-only structures provide weaker cues for stable multi-view reconstruction.
These limitations suggest promising directions for future work, including stronger segmentation priors, more robust cross-view correspondence strategies, scene-level vector generation, and topology-aware handling of open-path SVGs.

### v2

@@
 First, the rasterized input is lifted to a 3D representation and rendered under target camera poses, producing multi-view images of the object.
-Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining.
-Finally, during the raster to vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics.
+Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part-level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining.
+Finally, during the raster-to-vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics.
 The resulting SVGs exhibit strong geometric and color consistency across views, significantly reduce redundant paths, and retain fine structural details.
@@
 Editable vector graphics are central to modern design workflows, providing resolution independence, structural clarity, and precise editability.
-However, generating multi view consistent and fully editable vector graphics for a given object remains an open challenge.
+However, generating multi-view consistent and fully editable vector graphics for a given object remains an open challenge.
 When designers attempt to present an existing vector illustration from a new viewpoint, they often need to manually redraw or adjust its structure to approximate the desired perspective.
 This manual process is both time consuming and prone to geometric and stylistic inconsistencies.
-Adobe Illustrator’s Turntable (Beta) [1] is one of the few publicly available end-to-end tools that generate multiple views from a single 2D vector artwork, yet it provides constrained angle control and discloses no technical specification.
+Adobe Illustrator’s Turntable (Beta) [adobeTurntableHelp] is one of the few publicly available end-to-end tools that can generate multiple views from a single 2D vector artwork.
+However, it offers limited control over viewing angles and discloses no technical specifications.
 The task of automatically generating multi-view consistent and fully editable SVGs, however, has not been extensively explored.
-We define this task as synthesizing a set of path level editable SVGs that preserve geometric and stylistic coherence across viewpoints.
+We define this task as synthesizing a set of path-level editable SVGs that preserve geometric and stylistic coherence across viewpoints.
 Solving this problem is particularly valuable for creative workflows such as turntable style visualization, multi-view logo and icon generation, and geometry consistent SVG asset editing.
 Bridging the gap between generative modeling and structured vector representation is therefore critical for enabling scalable and reliable design pipelines.
-Recent advances in generative AI, particularly diffusion models[10, 29, 25, 30, 26, 46, 23], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [20, 33, 21, 38].
-Yet, these outputs remain raster based and lack the structured semantics and fine grained editability inherent to vector graphics.
-Meanwhile, existing vectorization approaches such as DiffVG [18], DeepSVG [3], and Im2Vec [28] focus on single view reconstruction and fail to maintain geometric consistency across views.
-This gap motivates a core objective: enabling generative models to produce multi view coherent and structurally editable representations directly in the SVG domain.
+Recent advances in generative AI, particularly diffusion models [ho2020ddpm, rombach2022ldm, nichol2021glide, saharia2022imagen, ramesh2022dalle2, zhang2023controlnet, mou2023t2iadapter], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [liu2023zero123, shi2023zero123pp, liu2023syncdreamer, voleti2024sv3d].
+Yet, these outputs remain raster-based and lack the structured semantics and fine grained editability inherent to vector graphics.
+Meanwhile, existing vectorization approaches such as DiffVG [li2020diffvg], DeepSVG [carlier2020deepsvg], and Im2Vec [reddy2021im2vec] focus on single-view reconstruction and fail to maintain geometric consistency across views.
+This gap motivates a core objective: enabling generative models to produce multi-view coherent and structurally editable representations in the SVG domain.
 To realize this objective, we propose a unified generative to vector framework that produces multi-view, geometrically consistent, fully editable SVGs from a single input.
 Our method consists of three stages.
-It first performs three dimensional aware multi-view generation to obtain geometry plausible novel views.
+It first performs 3D-aware multi-view generation to obtain geometrically plausible novel views.
 Then it performs cross-view refinement through spatially aligned segmentation and appearance harmonization, enhancing the consistency of object contours and colors across views in the SVG style domain.
 Finally, leveraging the multi-view consistent segmentation results, we perform raster-to-vector conversion followed by path consolidation and color correction, yielding compact, coherent, and easily editable multi-view SVGs.
-On multi-view SVG generation, our method produces noticeably cleaner and more stable results than existing vectorization tools.
-Compared with Adobe Illustrator’s Turntable, the path counts of our generated views stay much closer to the structure of the input SVG, reducing the path-count deviation from the input SVG by 26.5%.
-The variation in color usage between adjacent viewpoints is also reduced by 83.5%, indicating a far more consistent preservation of the input color design across rotations.
-In addition, our method lowers the average number of paths by 11.6%, resulting in lighter and more editable multi-view SVG sequences.
-The generated views exhibit coherent appearance transitions and show strong potential for creative design workflows, turntable-style visualization, and semantically aligned vector editing.
-Our contributions are summarized as follows:
+On multi-view SVG generation, our method produces noticeably cleaner and more stable results than a strong industrial baseline.
+Compared with Adobe Illustrator’s Turntable, our path count deviation ($\mathrm{RMSE}_{\text{path}}$) is about 23% lower, and the adjacent-view color count variation ($\overline{\Delta N}_{\text{color,nbr}}$) is about 86% lower.
+We also generate more compact SVGs, with about 7% fewer paths and about 17% fewer colors on average.
+In addition, our method yields smoother transitions on the yaw-pitch grid, with around 3% higher DINO feature consistency and around 2.5% higher DreamSim perceptual similarity on pitch changes [oquab2023dinov2, fu2023dreamsim].
+Finally, our vectorization module achieves high single-view fidelity measured by SSIM, LPIPS, and $\Delta E_{2000}$ [wang2004image, zhang2018unreasonable, sharma2005ciede2000].
+On our single-view benchmark, it attains the lowest $\Delta E_{2000}$ among the compared raster-to-SVG baselines while maintaining a compact topology [yue2024live, wang2024starvector, weber2001autotrace, jain2022vectorfusion].
 Unified Generative to Vector Framework.
-We introduce the first end-to-end framework to directly synthesize multi view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
-Spatially-Aware Cross-View Segmentation Refinement.
-We design a spatially-aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
+We introduce an end-to-end framework to directly synthesize multi-view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
+Spatially Aware Cross-View Segmentation Refinement.
+We design a spatially aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
 Structure Preserving Vector Consolidation.
@@
 2 Related Work
-2.1 Single image to multi view generation
-Turning one reference view into a usable turntable sequence has become a key precursor to our goal of editable multi-view SVGs. Existing approaches fall into two major categories:
-Diffusion-based multi-view image synthesis builds on high-fidelity 2D diffusion priors.
-These methods condition the denoising trajectory on target camera poses to synthesize each view directly in the image domain.
-Representative works include Zero-1-to-3 [20] and Zero123++ [33], where the former targets novel-view images rather than explicit 3D, and the latter improves cross-view consistency through better conditioning and training.
-Further, SyncDreamer [21] explicitly couples views during sampling via synchronized diffusion to align geometry and color, and SV3D [38] leverages video-latent diffusion to generate orbital view sequences from a single image.
-Despite strong appearance fidelity, residual independence across views often leads to geometric and color drift on thin structures and self occlusions.
-3D reconstruction and generation from single images.
-These methods predict a 3D asset from one or a few views and render all novel viewpoints from the shared geometry, which naturally improves multi-view geometric consistency over 2D diffusion synthesis.
-Representative approaches include mesh generators such as GET3D, PolyGen, MeshDiffusion [8, 24, 22], single-image reconstruction methods like LRM, TripoSR, UniQue3D [11, 36, 40], and latent 3D models including Shape-E and Trellis [14, 41].
-However, their outputs remain rasterized images without explicit vector paths or cross-view path correspondences, making them unsuitable for SVG-oriented workflows.
+2.1 Single-Image to 3D Vector Sketches
+Such methods predict sparse 3D contour curves from single-view renders [choi2024_3doodle, li2025_empowering, wang2025_viewcraft3d].
+Lacking filled regions and explicit colors, they inevitably lose the original SVG style.
+Instead, we desire to generate multi-view 2D SVGs that strictly preserve region geometry and design semantics.
+2.2 Single-image to multi-view generation
+Generating a turntable-like view sequence from a single reference image is a common precursor to editable multi-view SVGs. Existing approaches can be grouped into three categories.
+Diffusion-based multi-view image synthesis.
+These methods condition 2D diffusion models on target camera poses to synthesize views in the image domain.
+Representative works include Zero-1-to-3 and Zero123++ [liu2023zero123, shi2023zero123pp], SyncDreamer [liu2023syncdreamer], and SV3D [voleti2024sv3d].
+Despite strong appearance fidelity, view-wise sampling often causes geometric and color drift on thin structures and self-occlusions.
+Single-image 3D reconstruction and generation.
+These methods infer a shared 3D asset and render novel views from the same geometry, improving multi-view consistency over 2D diffusion.
+Representative approaches include mesh generators (GET3D, PolyGen, MeshDiffusion) [gao2022get3d, polygen, meshdiffusion], single-image reconstruction (LRM, TripoSR, UniQue3D) [hong2023lrm, triposr, unique3d], and latent 3D models (Shape-E, Trellis) [shapee, trellis2024].
+However, the outputs are still raster images without explicit vector paths or cross-view path correspondences.
 3D Gaussian Splatting (3DGS) as a shared representation.
-3DGS [15] represents a scene with anisotropic Gaussians and renders by fast differentiable splatting, so all views reproject the same underlying 3D field.
-This often yields strong cross view geometric stability and efficient multi view rendering.
-Recent works combine diffusion priors with 3DGS for rapid 3D asset creation [35, 7], and single image variants enable ultra fast inference [34].
-Although 3DGS provides efficient and geometrically consistent multi-view rendering, its outputs remain raster-based and do not explicitly expose vector path topology.
-Consequently, converting rendered contours into clean, closed, and editable paths requires an additional vectorization stage.
-2.2 Raster to Vector Generation
-Transforming multi-view consistent rasters into structured vector representations remains challenging because recovering continuous topology and geometry from discrete pixels is inherently ill-posed.
-Classical contour tracing methods such as Potrace and AutoTrace detect boundaries and fit Bézier curves or polygons.
-They perform well on icons or low-color imagery, but on smooth gradients and complex textures they often produce jagged, fragmented, or redundant paths.
-Learning-based approaches model vector paths parametrically.
-DiffVG [18] introduces a differentiable renderer for gradient-based optimization, and DeepSVG [3] and Im2Vec [28] leverage Transformer or VAE designs to improve geometric continuity and structural controllability.
-LIVE [45] decomposes images into hierarchical layers to preserve global topology, yet can over-segment fine details.
-More recent methods such as VectorFusion [13] and StarVector [39] incorporate diffusion priors or region fusion to obtain smoother and more coherent paths.
-Despite these advances, all vectorization methods still operate directly on pixel-level color transitions, which makes complex objects prone to fragmented or redundant paths.
-Introducing semantic segmentation before vectorization greatly simplifies the problem by isolating coherent part regions, enabling cleaner per-part conversion and more stable geometric structures.
-This highlights the need for a semantic layer, especially when extending vectorization to multi-view settings.
-2.3 Segmentation and Cross View Consistency
-While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation raises an additional requirement: achieving more consistent part assignments across viewpoints greatly improves cross-view coherence and visual stability.
-Yet conventional segmentation relies mainly on per-image pixel cues, making such consistency difficult; a spatially stable segmentation step is therefore needed before vectorization to provide more reliable part structures across views.
-Early two-dimensional methods [9, 4, 42, 17] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection.
-Foundation-scale models such as SAM, HQ-SAM, and SAM2 [16, 5, 27] provide accurate boundaries and promptable interaction, thus serving as strong building blocks for multi-view scenarios.
-However, their propagation mechanisms are primarily temporal rather than spatial, which limits their ability to maintain consistent part correspondences across viewpoints in static 3D scenes.
-Inspired by video object segmentation, memory-based propagation [6, 43] maintains features across frames to keep masks consistent over time.
-By analogy, multi-view consistency can be formulated as a spatial sequence problem on the viewing sphere, where each camera view is treated as a pseudo-temporal step and neighboring views act as spatial references.
-Existing work either stabilizes masks in raster space or improves single-view vectorization quality, but a bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored.
-This gap motivates our approach, which first achieves part-aware, cross-view consistent segmentation and then leverages these semantic units to drive compact and editable vector representations across views.
+3DGS [kerbl2023gaussians] enables efficient rendering by reprojecting a shared 3D field, often yielding stable multi-view geometry.
+Recent works combine diffusion priors with 3DGS for rapid asset creation [tang2023dreamgaussian, gaussiandreamer2023] and extend to single-image variants [splatterimage2024].
+Nevertheless, the results remain raster-based and require an additional vectorization stage to obtain clean, closed, and editable SVG paths.
+2.3 Raster-to-Vector Generation
+Converting rendered rasters into structured, editable vectors is challenging, especially when consistency must be maintained across multiple views.
+Classical tracing methods such as Potrace [selinger2003potrace] and AutoTrace [weber2001autotrace] work well for icons and low-color graphics, but often produce fragmented or redundant paths under smooth shading and complex textures.
+Prior work includes optimization-based methods, e.g., DiffVG [li2020diffvg] which optimizes control points via differentiable rendering, and learning-based methods such as DeepSVG [carlier2020deepsvg] and Im2Vec [reddy2021im2vec] built on Transformer [vaswani2017attention] or VAE [kingma2014vae] designs.
+For raster-to-SVG conversion, LIVE [yue2024live] preserves global topology via hierarchical layers but can over-segment fine structures, while VectorFusion [jain2022vectorfusion] and StarVector [wang2024starvector] incorporate diffusion priors or region-aware fusion for smoother paths.
+Despite these advances, many pipelines still rely heavily on local pixel-level color transitions, making complex objects prone to over-fragmentation in multi-view settings.
+Introducing part-level segmentation before vectorization isolates coherent regions, enabling cleaner per-part conversion and more stable structure and palette behavior across views.
+Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis.
+We then apply FLUX.1 dev [blackforestlabs2024flux] with a public vector-style LoRA [hu2021lora] adapter (SimpleVectorFlux [renderartist2024simplevectorflux]) to harmonize their appearance.
+The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere.
+During segmentation propagation, the Spatial Memory Selector retrieves the viewpoint-wise nearest memory entries.
+For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints.
+The resulting masks are vectorized by VTracer [vtracer2023] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
+2.4 Segmentation and Cross-view Consistency
+While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation additionally requires consistent part assignments across viewpoints.
+However, conventional segmentation relies on per-image pixel cues, making such consistency difficult; we therefore need a spatially stable segmentation step before vectorization to provide reliable part structures across views.
+Early two-dimensional methods [he2017maskrcnn, cheng2022mask2former, xie2021segformer, li2023maskdino] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection.
+Foundation-scale models such as SAM, HQ-SAM, and SAM2 [kirillov2023sam, cheng2023hqsam, ravi2024sam2] provide accurate boundaries and promptable interaction, but their propagation mechanisms are primarily temporal rather than spatial, which limits consistent part correspondences across viewpoints in static 3D scenes.
+Memory-based propagation in video object segmentation [cheng2021stcn, yang2023deAOT] maintains features across frames to keep masks consistent over time.
+In multi-view settings, we can similarly view cameras as a spatial sequence on the viewing sphere, treating each view as a pseudo-temporal step and using neighboring views as spatial references.
+A bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored.
+Our approach addresses this gap by first achieving part-aware, cross-view consistent segmentation, and then leveraging these semantic units to drive compact and editable vector representations across views.
 3 Method
-Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis.
-A lightweight LoRA [12] tuned FLUX [2] model is then applied to harmonize their appearance.
-The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere.
-During segmentation propagation, the Spatial Memory Selector retrieves the geometrically most relevant memory entries.
-For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints.
-The resulting masks are vectorized by VTracer [37] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
 As illustrated in Figure 1, our SVG360 framework begins with a 3D-aware rasterization stage (§3.1) that lifts a single-view SVG into geometrically plausible multi-view rasters.
@@
 To address this, we adopt a generative 3D-based approach to synthesize multi-view images.
-Although several recent single-image 3D methods (e.g., Sparc3D [19] and Hi3DGen [44]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs.
+Although several recent single-image 3D methods (e.g., Sparc3D [sparc3d2025] and Hi3DGen [hi3dgen2025]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs.
 Such geometry-only representations are unsuitable for our color consistent raster to SVG multi-view workflow.
-Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [41] as the backbone 3D generator.
+Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [trellis2024] as the backbone 3D generator.
 Trellis offers a balanced trade-off between geometric accuracy and texture consistency, and remains efficient and practical for producing coherent multi-view rasters from a single SVG input.
-Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [41].
-Its mesh decoder builds on FlexiCubes [32], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface.
+Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [trellis2024].
+Its mesh decoder builds on FlexiCubes [flexicubes2023], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface.
 While both decoders yield high-fidelity results, converting the latent to a mesh and then rasterizing from many viewpoints introduces extra discretization and baking steps, which may accumulate appearance deviations relative to the input SVG style and incur nontrivial runtime overhead.
@@
 While a 3D proxy preserves global geometric consistency across viewpoints, the reconstruction can still exhibit local errors, especially along edges and curved surfaces.
-View dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion.
-We apply appearance harmonization using FLUX.1-dev [2] fine-tuned with a LoRA [12] trained on flat-design SVG data.
+View-dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion.
+We apply appearance harmonization using FLUX.1-dev [blackforestlabs2024flux] with a public vector-style LoRA adapter (SimpleVectorFlux [renderartist2024simplevectorflux]).
 The harmonized renders increase intra-region homogeneity and boundary contrast, suppress spurious shading edges, and reduce color variation across views.
-The resulting rasters are both view-consistent and faithful to SVG-style appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.
+The resulting rasters are both view-consistent and faithful to a flat vector appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.
 3.2 Spatially Aligned Segmentation Propagation
@@
 To obtain cleaner and semantically stable vector paths in the subsequent vectorization stage, we need to maintain consistency of part-level identity across views.
-Relying on single view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence.
+Relying on single-view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence.
 Multi-view SVG generation inherently requires segmentation to be continuous in space.
@@
 Each target view is segmented under guidance from its nearest spatial neighbors, thereby maintaining part identity and boundary consistency at a global level.
-Spatially Sequential Scheduling.
+Spatially Sequential Scheduling
 We uniformly sample camera viewpoints on a unit sphere to obtain a globally distributed set of directions covering the entire view space.
@@
 To further enhance traversal smoothness, we apply a lightweight two-segment swap (2-opt) optimization to remove local discontinuities and maintain continuous transitions between neighboring viewpoints.
-Reference View Selection.
+Spatial Memory Selector
 After determining the traversal order, we construct a reference set for each target view to reuse information from geometrically adjacent and previously processed views.
@@
 We empirically set $k=6$ and $\tau=75^{\circ}$ in all experiments.
-This localized reference selection, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
+This Spatial Memory Selector mechanism, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
 Key Frame Initialization and Filtering.
 Segmentation starts from the key view $\theta_{0}$, which corresponds to the input SVG.
-After obtaining initial masks from the automatic segmentation model, we apply a lightweight post-processing to reduce redundancy and improve stability for subsequent propagation.
-Small regions smaller than either 200 pixels or 0.05% of the image area (whichever is larger) are removed, a single morphological closing with an elliptical kernel sized to about 0.5% of the shorter image dimension is used to smooth boundaries, and overlapping masks are suppressed using non-maximum suppression with an IoU threshold of 0.5.
-The cleaned masks serve as prompt inputs to the spatially SAM2, providing stable guidance for the following multi-view segmentation process.
-Residual Discovery.
+After obtaining initial masks from the automatic segmentation model, we apply lightweight postprocessing to reduce redundancy and improve stability for subsequent propagation.
+We remove regions whose area is smaller than the larger of 100 pixels and 0.05% of the image area.
+We then apply a single morphological closing with an elliptical kernel whose size is about 0.5% of the shorter image dimension to smooth boundaries.
+Finally, we suppress overlapping masks using non-maximum suppression with an IoU threshold of 0.5.
+The cleaned masks serve as prompts for Spatial SAM2, providing stable guidance for the subsequent multi-view segmentation process.
+Residual Discovery
 Since the initial segmentation relies solely on the key view $\theta_{0}$, previously unseen parts that appear in other viewpoints may remain undetected.
-After Spatial-SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions.
+After Spatial SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions.
 For each processed view $\theta_{t}$, we compute the uncovered foreground ratio $r$ based on the union of all current masks.
@@
         Merge new masks into existing results;
-        Propagate updates to neighboring views via Spatial-SAM2;
+        Propagate updates to neighboring views via Spatial SAM2;
  $k\leftarrow k+1$;
@@
 3.3 Vector Consolidation
-Figure 2: Qualitative comparison.
-The figure summarizes representative issues observed in Adobe Turntable: (a) geometric inconsistencies across adjacent views; (b) cluttered structures arising from overlapping thin components; (c) merging of parts with similar colors; (d) missing regions in certain viewpoints; (e) color drift and gradual loss of small details across views.
-Our method produces multi-view SVGs with stable geometry, clear part separation, and consistent color appearance.
-After segmentation, we crop each part into a transparent raster and feed it to VTracer [37], which converts pixel regions into path sets by color layering and spline fitting.
+After segmentation, we crop each part into a transparent raster and feed it to VTracer [vtracer2023], which converts pixel regions into path sets by color layering and spline fitting.
 Processing per part rather than per image reduces cross-part interference and yields a more compact path topology.
@@
 This consolidation reduces the number of paths and colors, improves topological simplicity and editability, and yields more consistent boundaries across neighboring views.
-To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric [31].
+To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric $\Delta E_{2000}$ [sharma2005ciede2000].
 We add a near-black bias to stabilize dark tones: colors in low lightness and low chroma preferentially map to black, otherwise to the closest non-black reference color.
@@
 For the 3D generation stage, we employ the TRELLIS-image-large model to generate multi-view rasterizations from the input SVG.
-To harmonize the rendered appearance, lightweight LoRA adapter is trained to adapt FLUX.1-dev to SVG-style appearance, following standard PEFT settings (rank=32).
-The text encoder is frozen, and the adapter is trained for 1,000 steps.
+To harmonize the rendered appearance, we use FLUX.1-dev [blackforestlabs2024flux] with a public vector-style LoRA adapter (SimpleVectorFlux [renderartist2024simplevectorflux, hu2021lora]), without any additional training.
+We follow the adapter’s recommended trigger tokens (v3ct0r, vector) and keep a fixed prompt template across all views for consistent style.
 All experiments are conducted on a single NVIDIA A100-80GB PCIe GPU.
 4.2 Evaluation Metrics
-We use two metrics to capture cross-view stability.
-$\mathrm{RMSE}_{\text{path}}$ measures how much each view’s path count deviates from the input SVG.
-$\overline{\Delta N}_{\text{color,nbr}}$ measures the change in color count between adjacent views, lower indicates more stable color behavior.
+To comprehensively evaluate our framework, we group our metrics into three aspects.
+SVG-Centric Statistics: We evaluate the compactness of generated SVGs using the average number of paths ($\overline{N}_{\text{path}}$) and colors ($\overline{N}_{\text{color}}$).
+To measure structural and appearance stability relative to the input, we use $\mathrm{RMSE}_{\text{path}}$ for path count deviation and $\overline{\Delta N}_{\text{color,nbr}}$ alongside the CIEDE2000 metric $\Delta E_{2000}$ [sharma2005ciede2000] to capture color count changes and perceptual color drift across views.
+Multi-View Continuity: For appearance continuity on the yaw-pitch viewing grid, we compute DINOv2 [oquab2023dinov2] feature consistency and DreamSim [fu2023dreamsim] perceptual similarity on adjacent view pairs.
+Specifically, we form yaw-adjacent pairs by grouping views with the same pitch and consecutive yaw steps, and pitch-adjacent pairs by grouping views with the same yaw and consecutive pitch levels.
+We report Yaw and Pitch by averaging scores over the corresponding adjacent pairs, and Overall by averaging over all adjacent pairs on the grid.
+Single-View Fidelity: We use standard image quality metrics, SSIM [wang2004image] and LPIPS [zhang2018unreasonable], to evaluate the fundamental raster-to-SVG reconstruction quality.
+Detailed mathematical formulations and computation specifics for all aforementioned metrics are provided in the supplementary material.
 4.3 Comparison
-Baseline Adobe Illustrator Turntable (Beta) is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork.
+Baselines.
+For multi-view generation, Adobe Illustrator’s Turntable (Beta) [adobeTurntableHelp] serves as our primary industrial baseline.
+It is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork.
 Its internal implementation, view grid, and technical specifications are not disclosed, and no programmatic interface is provided.
-We therefore use Turntable as the industrial baseline for our full pipeline comparison.
 According to the official documentation, Turntable supports approximately $\pm 120^{\circ}$ horizontal rotation in yaw (sampled every $15^{\circ}$) and three pitch levels at $\{-45^{\circ},0^{\circ},+45^{\circ}\}$.
-For fair comparison, we adopt the same camera poses defined by the Turntable configuration in our experiments.
-Because Turntable only allows manual per-view SVG export, we evaluate this baseline on 10 representative SVG examples from which complete view sequences can be consistently obtained.
-Quantitative Comparison Table 1 reports the numerical statistics for Adobe Turntable and our method.
-Our approach produces fewer paths and colors per view, indicating cleaner vector topology and a more compact palette.
-For cross-view stability, our method achieves lower $\mathrm{RMSE}_{\text{path}}$ and a substantially lower $\overline{\Delta N}_{\text{color,nbr}}$, showing that the generated views remain closer to the original SVG structure and maintain more consistent color behavior across viewpoints.
-Together, these results confirm that our multi-view SVGs are structurally simpler and exhibit smoother appearance transitions.
-Table 1: Comparison of multi-view SVG statistics between Adobe Turntable and our method.
-Metric Adobe Turntable Ours
-$\overline{N}_{\text{path}}$ 29.59 26.15
-$\overline{N}_{\text{color}}$ 4.80 3.37
-$\mathrm{RMSE}_{\text{path}}$  ( $\downarrow$ ) 16.37 12.03
-$\overline{\Delta N}_{\text{color,nbr}}$  ( $\downarrow$ ) 2.12 0.35
+For a fair comparison, we adopt the exact same camera poses defined by the Turntable configuration in our experiments.
+Since Adobe Turntable lacks an API and necessitates manual per-view export, large-scale evaluation is not practical to scale.
+To ensure rigorous comparison, we curated a representative benchmark of 30 SVGs across categories like vehicles, buildings, and animals.
+By fully exporting the 51-view yaw-pitch grid for each, we obtained a dense set of 1530 multi-view SVG outputs for strict quantitative and qualitative evaluation.
+For single view vectorization, we additionally benchmark our vectorization module on a small set of 200 simple single object SVGs and compare against LIVE [yue2024live], StarVector [wang2024starvector], and AutoTrace [weber2001autotrace].
+Figure 2: Qualitative comparison with Adobe Turntable.
+Target camera poses (yaw and pitch) are annotated at the top; !
+marks views that deviate from the target pose.
+(a) Car: unrealistic structural changes (e.g., incomplete or misplaced headlights) and inconsistent foreshortening/scale (e.g., wheels vs.
+chassis).
+(b) Torii gate: slender hanging components suffer from path crowding and tangling in certain views.
+(c) Scooter: distorted handlebar proportions in oblique views and structural inconsistency around the steering tube (head-tube region).
+(d) Turtle: missing small parts (e.g., nose) and locally incomplete shell regions in challenging views.
+Quantitative Comparison.
+Multi-View SVG Statistics.
+Table 1 compares the intrinsic properties of generated SVGs.
+Our approach produces highly compact vectors (avg.
+20.77 paths, 3.65 colors) while demonstrating exceptional stability across viewpoints.
+Compared to Adobe Turntable, we reduce path count deviation ($\mathrm{RMSE}_{\text{path}}$) from 14.22 to 10.95.
+Color stability improvements are particularly striking: we reduce adjacent-view color count variation ($\overline{\Delta N}_{\text{color,nbr}}$) from 2.01 to 0.29, and perceptual color drift ($\Delta E_{2000}$) from 3.22 to 1.19, substantially reducing the structural flickering and color inconsistencies common in the baseline.
+Table 1: SVG-centric multi-view comparison against Adobe Turntable.
+We report statistics of the generated multi-view SVGs, including compactness (fewer paths/colors indicate a more compact and editable SVG) and cross-view stability derived from SVG structure and palette behavior.
+Lower $\mathrm{RMSE}_{\text{path}}$, $\overline{\Delta N}_{\text{color,nbr}}$, and $\Delta E_{2000}$ indicate more stable structure and color across views.
+Method Compactness SVG structure Color consistency
+$\overline{N}_{\text{path}}$ $\overline{N}_{\text{color}}$ $\mathrm{RMSE}_{\text{path}}\downarrow$ $\overline{\Delta N}_{\text{color,nbr}}\downarrow$ $\Delta E_{2000}\downarrow$
+Adobe Turntable 22.34 4.41 14.22 2.01 3.22
+Ours 20.77 3.65 10.95 0.29 1.19
+Multi-View Continuity.
+Table 2 evaluates spatial continuity across the yaw-pitch viewing grid.
+Overall, our approach yields more coherent view-to-view transitions than Adobe Turntable.
+While Turntable attains competitive perceptual scores on simple yaw rotations, our method is more robust on the full grid, achieving higher overall DINO consistency (0.8864 vs.
+0.8611).
+Overall DreamSim is comparable (0.8392 vs.
+0.8388), but the advantage becomes clear under challenging pitch changes, where our method achieves higher Pitch DreamSim (0.7748 vs.
+0.7563), indicating smoother appearance transitions across viewpoints.
+Table 2: Multi view continuity comparison against Adobe Turntable on a yaw and pitch camera grid.
+We measure local transition smoothness between adjacent views using DINOv2 feature consistency and DreamSim perceptual similarity, reporting yaw, pitch, and overall scores.
+Method DINO Consistency DreamSim Similarity
+[…]
+Adobe Turntable 0.9005 0.8053 0.8611 0.8973 0.7563 0.8388
+Ours 0.9074 0.8567 0.8864 0.8848 0.7748 0.8392
+Single-View Raster-to-SVG Fidelity.
+Table 3 validates our vectorization module against state-of-the-art models.
+Our method yields the highest visual fidelity (SSIM: 0.926, LPIPS: 0.020, $\Delta E_{2000}$: 0.408) while striking an optimal balance in structural compactness.
+Specifically, while StarVector produces slightly fewer paths (15.2) at the cost of poor fidelity ($\Delta E_{2000}$: 17.706), and AutoTrace yields reasonable similarity (SSIM: 0.812) but heavily fragmented topology (448.2 paths), our method avoids these trade-offs, delivering near-perfect reconstruction with only 19.75 paths.
+Table 3: Single-view (initial input view) raster-to-SVG comparison. We render predicted SVGs back to rasters and evaluate against the input raster.
+[…]
+LIVE [yue2024live] 0.803 0.065 7.364 18.2
+StarVector [wang2024starvector] 0.697 0.153 17.706 15.2
+AutoTrace [weber2001autotrace] 0.812 0.063 2.792 448.2
+Ours 0.926 0.020 0.408 19.75
 Qualitative Comparison.
-Figure 2 summarizes the typical failure patterns we observe when applying Adobe Turntable to single-SVG multi-view generation.
-Across multiple objects and views, Adobe Turntable frequently exhibits geometric inconsistencies, structural confusion, unstable camera poses, view-dependent clipping, and noticeable color drift.
-In contrast, our method maintains coherent geometry, clear part separation, and consistent appearance across views.
-In terms of geometry, object structures do not remain consistent across views.
-In example (a), components that should be rigid, such as the roof rack and the headlights—change noticeably between yaw $0^{\circ}$ and $90^{\circ}$, with the rack switching from two crossbars to a flat plate and the paired headlights collapsing into a single central light.
-Examples (b) and (c) further show that when the input contains multiple thin or closely spaced parts, the generated views often introduce overlaps, clutter, or unintended merging, making individual components difficult to distinguish.
-Viewpoint-related inconsistencies are also apparent.
-In examples (a) and (e), the result at yaw $90^{\circ}$ deviates from a clean side view, exhibiting a slight tilt and geometric deformation.
-When the input SVG spans a large spatial extent, Turntable does not normalize its scale, causing portions of the object to be clipped outside the rendering canvas, as observed in both (a) and (e).
-Color stability across views can also be unreliable.
-In example (e), the sesame seeds on the burger change brightness across viewpoints, and those on the far side gradually diminish as the camera rotates.
-In example (c), regions with similar color are merged into a single path, reducing structural separation.
-In contrast, our method relies on explicit 3D representation, controlled camera poses, and consistent color and path processing.
-As a result, the generated multi-view SVGs maintain stable geometry, clear part separation, and coherent appearance across all viewpoints.
+Figure 2 provides a visual comparison with Adobe Turntable.
+Overall, the baseline often struggles to maintain cross-view 3D consistency for single-SVG multi-view generation.
+It is prone to geometric/foreshortening distortions, cluttered paths on thin or overlapping structures, and occasional pose drift.
+Under challenging viewpoints, it may also miss small semantic parts.
+In contrast, by leveraging an explicit 3D representation with strict pose control, our approach maintains more stable geometry, cleaner part separation, and coherent multi-view appearances.
 4.4 Ablation Study
-Figure 3: Segmentation comparison of Spatial-SAM2 and the original SAM2 tracking mode at view 47.
-The bottom row shows the each six reference views used for memory support, with view 0 serving as the initial key frame.
-Results shown here correspond to the first iteration before any subsequent refinement.
+Figure 3: Segmentation comparison between the original SAM2 tracking mode (left) and Spatial SAM2 (ours, right) at view 47.
+For each method, we show the six reference views used for memory support.
+Tracking SAM2 uses views 41-46, while Spatial SAM2 uses views 24-26, and 44-46.
+Results correspond to the first iteration before any subsequent refinement.
 Figure 3 illustrates the difference between the original Tracking SAM2 (Video Segmentation) and our Spatial SAM2 when propagating masks across multi-view settings.
@@
 As a result, even at spatially distant viewpoints, structures such as wheels remain stable and consistently segmented across views.
-Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation free VTracer baseline.
 As shown in Figure 4, we further compare the impact of four segmentation strategies on the final SVG outputs.
-Spatial-SAM2 produces the most stable and coherent results.
+Spatial SAM2 produces the most stable and coherent results.
 The tracking mode of SAM2 still exhibits occasional misses in certain viewpoints.
@@
 These findings confirm that enforcing spatially consistent segmentation is critical for producing high-quality multi-view SVGs.
+Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation-free VTracer baseline.
 5 Conclusion
@@
 These limitations suggest promising directions for future work, including stronger segmentation priors, more robust cross-view correspondence strategies, scene-level vector generation, and topology-aware handling of open-path SVGs.
+Beyond these extensions, we also see opportunities to better exploit vector-specific structure, e.g., enforcing cross-view consistency of palettes and layer hierarchies.
+Supplementary Material
+Appendix 0.A Implementation Details
+Our full system consists of three main stages: (1) single-SVG-conditioned multi-view image generation, (2) cross-view part-aware segmentation and propagation, and (3) single-view raster-to-SVG vectorization for each generated view.
+3D Generation and Harmonization.
+For the multi-view raster generation, we employ the TRELLIS-image-large model [trellis2024] as our 3D prior.
+To harmonize the rendered appearance and reduce view-dependent shading artifacts, we use FLUX.1-dev [blackforestlabs2024flux] combined with the SimpleVector Flux [renderartist2024simplevectorflux] LoRA [hu2021lora] adapter.
+We set the inference steps for FLUX to 20 with a guidance scale (CFG) of 3.5.
+To maintain a consistent SVG style across all viewpoints, we use a fixed prompt template: “A clean vector illustration of a [object class], flat colors, v3ct0r, vector style, white background”.
+Spatial SAM2 and Vector Consolidation.
+In the Spatial SAM2 module [ravi2024sam2], we set the neighborhood size for the Spatial Memory Selector to $k=6$ and the angular threshold to $\tau=75^{\circ}$.
+During the vector consolidation stage, regions with an area smaller than 100 pixels (or 0.05% of the image area, whichever is larger) are treated as noise and filtered out.
+To resolve stroke artifacts from anti-aliasing, we explicitly correct hue and brightness deviations by mapping source colors to the nearest reference palette entry using the CIEDE2000 metric ($\Delta E_{2000}$[sharma2005ciede2000].
+Appendix 0.B Additional Experiments and Comparisons
+0.B.1 Comparison with Multi-View Image Generators
+While our primary goal is multi-view SVG generation, the geometric stability of the intermediate multi-view rasters is crucial.
+In Figure S1, we qualitatively compare our 3D-guided rendering approach, which utilizes the TRELLIS-image-large [trellis2024] model as a 3D prior, against state-of-the-art multi-view image generators: Hunyuan3D [yang2024hunyuan3d], MV-Adapter [huang2024mvadapter], and Zero123++ [shi2023zero123pp].
+Figure S1: Qualitative comparison of multi-view image generation.
+Baselines exhibit strong view-dependent distortions and shading drift, whereas our method (based on TRELLIS [trellis2024]) maintains coherent structure and stable appearance.
+0.B.2 Qualitative Vectorization Comparison
+We provide qualitative comparisons corresponding to the single-view raster-to-SVG fidelity evaluation (Table 3 in the main text).
+Figure S2 presents the vectorization results of identical input raster images.
+Recent learning-based baselines exhibit distinct artifacts: LIVE [yue2024live] produces blurry edges and color bleeding, while StarVector [wang2024starvector] merges neighboring regions excessively, losing local contrast.
+Classical methods like AutoTrace [weber2001autotrace] fragment smooth color fields into thousands of polygons, resulting in noisy outlines.
+In comparison, our vectorization module, which leverages part-aware segmentation and topological consolidation, delivers clean, visually consistent SVGs with accurate color boundaries and minimal redundancy.
+Figure S2: Qualitative vectorization results under identical viewpoints.
+Our method delivers clean SVGs with accurate color boundaries, whereas baselines suffer from blurry edges, color bleeding, or severe polygon fragmentation.
+Appendix 0.C Analysis of Key Modules
+0.C.1 Visual Analysis of Appearance Harmonization
+We first examine the role of appearance harmonization in our pipeline.
+As shown in Figure S3, raw 3D renderings often exhibit view-dependent shading variations, local intensity gradients, and mild chromatic inconsistencies.
+Although these artifacts do not alter the underlying geometry, they introduce appearance noise that can interfere with downstream segmentation and raster-to-SVG conversion.
+After refinement with FLUX.1-dev [blackforestlabs2024flux] and the public vector-style LoRA adapter (SimpleVectorFlux [renderartist2024simplevectorflux, hu2021lora]), the rendered views become closer to a flat vector-style appearance while preserving the original object structure.
+The difference maps are concentrated primarily around shading regions, and the consistently high SSIM scores indicate that the refinement mainly affects appearance rather than geometry.
+In practice, this harmonization improves intra-region uniformity, sharpens region boundaries, and suppresses spurious shading edges, thereby making subsequent segmentation more reliable and the following vector consolidation more stable across viewpoints.
+Figure S3: Comparison between raw 3D renderings (left), refined results using fine-tuned FLUX-LoRA (middle), and their difference maps with SSIM scores (right) across multiple viewpoints.
+The high SSIM scores and localized difference maps indicate that the harmonization primarily removes appearance-level shading artifacts while preserving the underlying object structure.
+0.C.2 Single-View Evidence for Segmentation-Guided Vector Consolidation
+We next examine the effect of our vectorization stage in the single-view setting.
+This analysis complements the multi-view results by showing that the same design also improves SVG reconstruction quality for individual views.
+To make this effect explicit, we compare our full pipeline with a direct raster-to-vector baseline.
+Specifically, we use direct VTracer as the baseline, where the entire image is vectorized globally without part-aware masks or explicit palette alignment.
+This provides a meaningful reference because VTracer is also the tracing backend used in our method.
+The comparison therefore highlights the benefit of the additional design choices introduced by our pipeline, rather than differences in the underlying tracer itself.
+As shown in the path-line visualizations in Figure S4, direct raster-to-vector conversion remains sensitive to anti-aliasing artifacts and sub-pixel color variation, often producing fragmented micro-paths, overlapping boundaries, and locally unstable topology.
+In contrast, our full pipeline yields cleaner path layouts and sharper region boundaries.
+These improvements arise from the combined effect of three design choices.
+First, part-aware segmentation decomposes the raster input into structurally meaningful regions before tracing.
+Second, region-wise vectorization reduces cross-part interference and prevents paths from crossing semantic boundaries.
+Third, palette alignment maps traced colors back to the reference SVG palette, suppressing residual chromatic noise and improving color faithfulness.
+Table S1 summarizes the quantitative comparison.
+Relative to direct VTracer, our full pipeline improves SSIM from 0.833 to 0.926, reduces LPIPS from 0.0206 to 0.0200, and substantially lowers perceptual color drift ($\Delta E_{2000}$) from 1.340 to 0.408, while also slightly reducing the average number of paths from 20.4 to 19.75.
+These results indicate that the proposed vector consolidation strategy contributes not only to multi-view consistency, but also to high-fidelity single-view SVG reconstruction.
+Figure S4: Comparison between direct VTracer vectorization and our full vectorization pipeline.
+By combining part-aware segmentation, region-wise tracing, and palette alignment, our method suppresses noise-driven micro-paths and redundant intersections, yielding cleaner topology and more accurate color boundaries.
+Table S1: Single-view comparison between direct VTracer and our full vectorization pipeline.
+The gains in SSIM, LPIPS, and $\Delta E_{2000}$ show that segmentation-guided tracing and palette alignment substantially improve SVG fidelity while preserving a compact vector structure.
+Method SSIM $\uparrow$ LPIPS $\downarrow$ $\bm{\Delta E_{2000}}$ $\downarrow$ NumPaths $\downarrow$
+VTracer (Direct) 0.833 0.0206 1.340 20.4
+Ours 0.926 0.0200 0.408 19.75
+Appendix 0.D Full 360 Multi-View SVG Sequences
+To comprehensively demonstrate the geometric stability and robust color consistency of SVG360, we present full 51-view turntable sequences for several representative objects in Figures S5 to S8.
+These uncurated grids show both the rendered SVGs and their underlying path topologies.
+Our method consistently avoids the structural flickering, shape collapse, and severe semantic drift that often plague purely 2D generative approaches or direct auto-tracing methods, maintaining high fidelity to the original SVG style across all viewing angles.
+Figure S5: Full multi-view sequence for “Blue Car”. Top rows show the rendered SVGs; bottom rows visualize the corresponding vector paths.
+Figure S6: Full multi-view sequence for “Torii Gate”. Top rows show the rendered SVGs; bottom rows visualize the corresponding vector paths.
+Figure S7: Full multi-view sequence for “Hamburger”. Top rows show the rendered SVGs; bottom rows visualize the corresponding vector paths.
+Figure S8: Full multi-view sequence for “Turtle”. Top rows show the rendered SVGs; bottom rows visualize the corresponding vector paths.
+Appendix 0.E Evaluation Metrics and Protocol
+Our evaluation is divided into three parts:
+  • SVG-centric statistics, which measure the compactness and cross-view stability of the generated SVGs.
+  • Multi-view continuity, which evaluates whether adjacent views preserve smooth feature and perceptual transitions on the yaw-pitch grid.
+  • Single-view raster-to-SVG fidelity, which measures how faithfully each generated SVG reproduces its raster target.
+These metrics are designed to jointly evaluate editability, structural consistency, color stability, and visual smoothness.
+0.E.1 Common evaluation protocol
+View set and notation.
+For each object, we evaluate a set of rendered viewpoints $\mathcal{V}$ that matches the Adobe Turntable [adobeTurntableHelp] configuration, with yaw sampled every $15^{\circ}$ within its supported range and pitch levels in $\{-45^{\circ},0^{\circ},+45^{\circ}\}$.
+For each view $v\in\mathcal{V}$, a method outputs an SVG $S_{v}$.
+We rasterize each SVG back into an image $I_{v}$ using identical rendering settings across methods, including resolution, background, and antialiasing.
+Adjacency on the yaw-pitch grid.
+We arrange all views into a 2D yaw-pitch grid.
+Let $\mathcal{E}_{\text{yaw}}$ denote the set of yaw-adjacent pairs with the same pitch and consecutive yaw steps.
+Let $\mathcal{E}_{\text{pitch}}$ denote the set of pitch-adjacent pairs with the same yaw and consecutive pitch levels.
+We define the full adjacent-pair set as
+$\mathcal{E}=\mathcal{E}_{\text{yaw}}\cup\mathcal{E}_{\text{pitch}}.$
+Dataset aggregation.
+For metrics computed per object, we first compute the metric for each object and then report the average across all objects.
+For metrics computed per view or per adjacent pair, we first average within each object and then average across objects.
+0.E.2 SVG-centric statistics
+These metrics characterize the compactness and stability of generated multi-view SVG outputs.
+Average number of paths $\overline{N}_{\text{path}}$.
+For each SVG $S_{v}$, we count the number of visible shape paths, excluding non-geometry elements such as defs, clipPath, and masks. Let $N_{\text{path}}(v)$ be the path count of view $v$. We define
+$\overline{N}_{\text{path}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}N_{\text{path}}(v).$
+This metric reflects vector complexity and practical editability. Fewer paths usually indicate a more compact SVG, although this should be interpreted together with fidelity metrics.
+Average number of colors $\overline{N}_{\text{color}}$.
+We extract all visible fill colors from the shape paths in $S_{v}$, represented in sRGB space. Let $N_{\text{color}}(v)$ be the number of unique fill colors in view $v$. We define
+$\overline{N}_{\text{color}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}N_{\text{color}}(v).$
+This metric measures palette compactness. A smaller number of colors often suggests reduced spurious variation and better downstream editability.
+Path-count stability $\mathrm{RMSE}_{\text{path}}$.
+Let $N_{\text{path}}^{\text{in}}$ be the number of visible shape paths in the input SVG. We compute the deviation of each generated view from the input complexity:
+$\mathrm{RMSE}_{\text{path}}=\sqrt{\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\left(N_{\text{path}}(v)-N_{\text{path}}^{\text{in}}\right)^{2}}.$
+This metric evaluates structural stability across views. A lower value indicates that the method preserves similar vector complexity under viewpoint changes.
+Neighbor color-count change $\overline{\Delta N}_{\text{color,nbr}}$.
+For each adjacent pair $(v_{i},v_{j})\in\mathcal{E}$, we compute the absolute difference in the number of unique colors:
+$\overline{\Delta N}_{\text{color,nbr}}=\frac{1}{|\mathcal{E}|}\sum_{(v_{i},v_{j})\in\mathcal{E}}\left|N_{\text{color}}(v_{i})-N_{\text{color}}(v_{j})\right|.$
+This metric measures palette stability across neighboring views. Lower values indicate smoother color behavior under small viewpoint changes.
+Perceptual color drift $\Delta E_{2000}$ on adjacent views.
+We use the CIEDE2000 distance [sharma2005ciede2000] to measure perceptual color differences between corresponding regions in adjacent views.
+To ensure geometric alignment, corresponding regions are matched via an IoU-based heuristic.
+For a matched region pair $(p,q)$, let $\mu_{v,p}$ denote the mean Lab color of region $p$ in view $v$.
+We define
+$\Delta E_{2000}=\frac{1}{|\mathcal{E}|}\sum_{(v_{i},v_{j})\in\mathcal{E}}\frac{1}{|\mathcal{M}_{v_{i}\rightarrow v_{j}}|}\sum_{(p,q)\in\mathcal{M}_{v_{i}\rightarrow v_{j}}}\Delta E_{00}\big(\mu_{v_{i},p},\mu_{v_{j},q}\big).$
+Here, $\mathcal{M}_{v_{i}\rightarrow v_{j}}$ denotes the set of matched region pairs between adjacent views $v_{i}$ and $v_{j}$.
+In practice, adjacent pairs with no valid region matches are excluded from the computation.
+This metric captures perceptual color consistency across neighboring views.
+Lower values indicate less viewpoint-dependent color drift.
+0.E.3 Multi-view continuity
+We evaluate whether neighboring views change smoothly on the yaw-pitch grid.
+Following the main paper, we report Yaw and Pitch scores by averaging over the corresponding adjacency sets, and Overall scores by averaging over all adjacent pairs on the full grid.
+DINO feature consistency.
+For an adjacent image pair $(I_{v_{i}},I_{v_{j}})$, we extract DINOv2 [oquab2023dinov2] patch-token features $\{f_{v_{i}}^{p}\}_{p=1}^{P}$ and $\{f_{v_{j}}^{q}\}_{q=1}^{P}$.
+For each patch $p$ in $I_{v_{i}}$, we find its nearest patch in $I_{v_{j}}$ using cosine similarity:
+$q^{\ast}(p)=\arg\max_{q}\ \frac{\langle f_{v_{i}}^{p},f_{v_{j}}^{q}\rangle}{\|f_{v_{i}}^{p}\|\|f_{v_{j}}^{q}\|}.$
+The adjacent-pair DINO score is defined as
+$S_{\text{DINO}}(I_{v_{i}},I_{v_{j}})=\frac{1}{P}\sum_{p=1}^{P}\frac{\langle f_{v_{i}}^{p},f_{v_{j}}^{q^{\ast}(p)}\rangle}{\|f_{v_{i}}^{p}\|\|f_{v_{j}}^{q^{\ast}(p)}\|}.$
+We then compute
+$\mathrm{DINO}_{\text{Yaw}}=\frac{1}{|\mathcal{E}_{\text{yaw}}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{\text{yaw}}}S_{\text{DINO}}(I_{v_{i}},I_{v_{j}}),$
+$\mathrm{DINO}_{\text{Pitch}}=\frac{1}{|\mathcal{E}_{\text{pitch}}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{\text{pitch}}}S_{\text{DINO}}(I_{v_{i}},I_{v_{j}}),$
+and
+$\mathrm{DINO}_{\text{Overall}}=\frac{1}{|\mathcal{E}|}\sum_{(v_{i},v_{j})\in\mathcal{E}}S_{\text{DINO}}(I_{v_{i}},I_{v_{j}}).$
+This metric captures feature-level structural continuity across adjacent views. Higher values indicate smoother transitions.
+DreamSim perceptual similarity.
+Using the same adjacency sets, we replace the pairwise similarity function with DreamSim [fu2023dreamsim]:
+$\mathrm{DreamSim}_{\text{Yaw}}=\frac{1}{|\mathcal{E}_{\text{yaw}}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{\text{yaw}}}\mathrm{DreamSim}(I_{v_{i}},I_{v_{j}}),$
+$\mathrm{DreamSim}_{\text{Pitch}}=\frac{1}{|\mathcal{E}_{\text{pitch}}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{\text{pitch}}}\mathrm{DreamSim}(I_{v_{i}},I_{v_{j}}),$
+and
+$\mathrm{DreamSim}_{\text{Overall}}=\frac{1}{|\mathcal{E}|}\sum_{(v_{i},v_{j})\in\mathcal{E}}\mathrm{DreamSim}(I_{v_{i}},I_{v_{j}}).$
+This metric captures perceptual continuity in a way that better aligns with human visual judgment. Higher values indicate smoother adjacent-view appearance.
+0.E.4 Single-view raster-to-SVG fidelity
+These metrics evaluate reconstruction quality in the single-view setting, independently of cross-view propagation.
+Rendering for evaluation.
+For each predicted SVG $S$, we rasterize it into an image $I_{\text{svg}}$ at the same resolution as the target raster image $I_{\text{in}}$, using identical background and antialiasing settings.
+All metrics below are computed between $I_{\text{svg}}$ and $I_{\text{in}}$, and the final reported score is the average over the evaluation set.
+SSIM.
+We compute the structural similarity index (SSIM) [wang2004image]:
+$\mathrm{SSIM}(I_{\text{svg}},I_{\text{in}}).$
+This metric measures structural fidelity and local contrast preservation. Higher values indicate better reconstruction quality.
+LPIPS.
+We compute LPIPS [zhang2018unreasonable] using the standard pretrained perceptual network:
+$\mathrm{LPIPS}(I_{\text{svg}},I_{\text{in}}).$
+This metric captures perceptual reconstruction quality beyond raw pixel matching. Lower values indicate better perceptual fidelity.
+Single-view $\Delta E_{2000}$.
+We compute the mean CIEDE2000 distance [sharma2005ciede2000] in Lab space between $I_{\text{svg}}$ and $I_{\text{in}}$, optionally restricted to a foreground region:
+$\Delta E_{2000}^{\text{sv}}=\frac{1}{|\Omega|}\sum_{x\in\Omega}\Delta E_{00}\big(\operatorname{Lab}(I_{\text{svg}}(x)),\operatorname{Lab}(I_{\text{in}}(x))\big),$
+where $\Omega$ denotes the evaluation region, either the full image or a foreground mask. This metric measures perceptually meaningful color fidelity. Lower values indicate better color reproduction.
+Path count (single view).
+For each single-view SVG output, we report the number of visible shape paths. This complements raster fidelity metrics by indicating vector complexity and editability.
+Appendix 0.F Additional Remarks
+The above metrics are designed to evaluate complementary aspects of multi-view SVG generation.
+SVG-centric statistics focus on compactness and stability of the vector representation.
+Multi-view continuity metrics evaluate smoothness across adjacent views on the yaw-pitch grid.
+Single-view SSIM, LPIPS, and $\Delta E_{2000}^{\text{sv}}$ measure the fidelity of the vectorized view itself.
+Together, these metrics provide a more complete assessment than either raster fidelity or vector compactness alone.